# Robust and Computation-Aware Gaussian Processes

### v1

@@ 7.2 High-Throughput Bayesian Optimization
We consider 4 tasks: Hartmann6D, Lunar12D, Rover60D and Lasso-DNA180D (Section H.2). Figure 3 reports the best value found during BO, averaged over 20 trials, with $\pm$1 standard deviation. To ensure fair comparisons despite outlier contamination, we report the uncontaminated mean best value. In the top row (columns 1–4), using EULBO-EI and under asymmetric outliers, RCaGP clearly outperforms all baselines on Lasso DNA 180D, and ranks first or on par with CaGP on Hartmann 6D, Lunar12D, and Rover60D. SVGP and RCSVGP display the same performance. The bottom row (columns 1-4) shows results with Determinantal Point Process (DPP)-based inducing point selection [30], still in the case of asymmetric outliers. RCaGP remains the top performer on DNA and performs similarly to CaGP on other tasks, with RCSVGP slightly ahead on Hartmann. Substituting EI with Thompson Sampling in DPP-BO leads to equal performance across baselines (Figure S1).
Next, Figure S2 presents results with TurBO, a BO variant tailored to high dimensions [8]. RCaGP exhibits strong average performance, and notably, it is the only method that never ranks last in a statistically significant manner. Lastly, as one might not know beforehand whether outliers will affect the trial, column 5 in Figure 3 shows results for EULBO EI on Hartmann and Lunar without outliers. RCaGP maintains a clear advantage on Hartmann, while all baselines perform similarly on Lunar.
Together, these results highlight RCaGP’s versatility and robustness across inducing point allocation strategies (EULBO, DPP), acquisition functions (EI, TS), BO behavior (global BO, trust-region BO), and noise regimes (with or without outliers). Its consistent superiority supports the central claim: combining robustness from RCGP with the uncertainty calibration of CaGP leads to a synergistic improvement—not just an additive one—across a wide range of BO challenges.
@@ 8 Discussion
  • Appendix A gives the results of additional experiments mentioned in the main text. These include experiments on UCI regression datasets with uniform and focused outliers (Table S1 and Table S2), BO results for DPP-BO-TS (Figure S1), TuRBO (Figure S2), and dataset-specific results for the ablation study conducted in Section 7.3, where $c$ is varied.
  • Appendix B gives the computational complexity of RCaGP.
  • Appendix C formally derives the ELBO for RCaGP introduced in Section 3.2.
  • Appendix D contains the proof for the robustness property presented in Proposition 4.1.
  • Appendix E proves Proposition 4.2, showing that RCaGP’s uncertainty estimates capture the worse-case error over all latent functions.
  • Appendix F contains additional theoretical results: a proposition showing the convergence in mean of RCaGP in RKHS norm (Section F.2), and a proposition showing the lack of robustness of CaGP (Section F.1).
  • Appendix G provides further details regarding the expert-driven prior mean introduced in Section 5 and employed during the ablation study in Section 7.3.
  • Appendix H provides additional details about the UCI regression datasets (Section H.1), the test functions employed in BO (Section H.2), and the outlier-contamination protocols studied in this work (Section H.3).
  • Appendix I provides tables that report the hyperparmeters used for UCI regression and high-throughput BO experiments (Section I.1) and the description of the computing resources used for our experiments (Section I.2).
@@ Appendix A Additional experiments
The results of the UCI regression experiments in the presence of uniform and focused outliers are presented in Table S1 and Table S2, respectively. The MAE and NLL metrics indicate that RCaGP consistently outperforms the baseline methods across most datasets. These findings demonstrate the robustness and versatility of RCaGP in handling different types of outliers. However, this performance improvement comes at the cost of increased computational time compared to the baselines.
The results of Bayesian optimization (BO) experiments using the Thompson Sampling (TS) acquisition function (AF) and Trust-Region BO (TuRBO) are presented in Figure S1 and Figure S2, respectively. In experiments utilizing the TF acquisition function, no model demonstrates a statistically significant performance advantage. Furthermore, the best values identified by these models are generally lower than those obtained with the BO-EI method. We hypothesize that this performance degradation arises from the reliance of the TS approach on posterior sampling, which may reduce the effectiveness of the robustness properties inherent in RCaGP. Subsequent BO experiments using the TuRBO framework show improved performance across all models and tasks—except for the Lunar task. In this case, TuRBO yields lower results for all models. This suggests that the presence of outliers in the Lunar task may impede TuRBO’s ability to effectively explore and identify the global optimum.
The MAE and NLL plots for varying values of $\epsilon$ in the weight function $w$ are shown in Figure S3. Across all datasets, RCaGP consistently yields lower MAE and NLL values than RCSVGP, indicating superior predictive performance. Furthermore, both metrics improve when $\epsilon$ is set to 0.1 or 0.15, but begin to deteriorate once $\epsilon$ exceeds 0.2. This trend suggests that aligning $\epsilon$ with the true proportion of outliers in the data enhances model robustness and overall performance.
Table S1: UCI Regression datasets results in the presence of uniform outliers with $\epsilon=0.1$. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits.
@@
Table S2: UCI Regression datasets results in the presence of focused outliers with $\epsilon=0.1$. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits.
@@ Appendix C Evidence Lower Bound for RCaGP
$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{k}_{i}({\bf{X}},{\bf{X}}){\bf{J}}_{\bf{w}}^{-1/2})-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}})^{\top}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}}))+\hat{\mu}_{i}({\bf{X}})^{\top}\nu$
@@
[…]
Next, we apply the identity ${\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}={\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})$ (see [2]):
@@ Appendix D Robustness Property
where we define $\hat{{\bf{G}}}={\bf{I}}-\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}})(\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}}))^{+}$ and + denotes the Moore-Penrose inverse.
@@
PIF of RCaGP We now prove Proposition 4.1: Let ${\bf{f}}\sim{\mathcal{GP}}(m,k)$ denote the RCaGP prior, and let $i\in{0,\dots,n}$ represent the number of actions in RCaGP. Assume the observation noise is $\varepsilon\sim{\mathcal{N}}(\mathbf{0},{\sigma_{\mathrm{noise}}^{2}}{\bf{I}})$. Define constants $C_{k}^{\prime}\in;k=1,2,3$, which are independent of $y_{m}^{c}$. For any $i$ and assuming $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$, the PIF of RCaGP is given by:
@@
[…]
Next, we derive the bound for the term (2). Following [2], we have that
@@
for a constant $q<+\infty$. The second equality comes from the definition of $\tilde{{\bf{v}}}_{i}$, and the fourth row follows the Cauchy-Schwarz inequality. Finally, the last inequality holds since $\tilde{{\bf{K}}}_{i}=(\tilde{{\bf{C}}}_{i}^{-1}-\tilde{{\boldsymbol{\Sigma}}}_{i})$. Applying results from [2], we obtain
@@
where $\bar{C}_{11}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{7}+\bar{C}_{8})^{2}$ and $\bar{C}_{12}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{9}+\bar{C}_{10})^{2}$. The terms $\bar{C}_{7},\bar{C}_{8},\bar{C}_{9},\bar{C}_{10}$ correspond to $\tilde{C}_{6},\tilde{C}_{8},\tilde{C}_{7},\tilde{C}_{9}$ in [2].
@@ Appendix E RCaGP Capture the Worst-case Errors
where ${\mathcal{D}}$ denotes the observation set. Then, applying [20, Lemma 3.9] provides
@@
since ${\bf{x}}\neq{\bf{x}}_{j}$, it holds that $k^{w}({\bf{x}},{\bf{X}})=k({\bf{x}},{\bf{X}})$. By definition, we have $k^{w}({\bf{X}},{\bf{X}})=\tilde{{\bf{K}}}={\bf{K}}+\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}{\bf{J}}_{\bf{w}}$, and following in [45, Eq. (S42)], it holds that $\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}=\tilde{{\bf{C}}_{i}}$. Therefore, we obtain
@@
$\quad\;k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})$ (S100)
@@
The proof follows immediately from Proposition 4.2 and notice that $h/\|h\|_{{\mathcal{H}}_{k^{w}}}$ has unit norm. Corollary E.1 implies that the computational uncertainty provides the pointwise bound relative to RCGP’s mathematical posterior, and the combined uncertainty is the pointwise bound relative to the shifted latent function.

### v2

@@ 7.3 High-Throughput Bayesian Optimization
We consider 4 tasks: Hartmann6D, Lunar12D, Rover60D and Lasso-DNA180D (Section H.2). Figure˜4 reports the best value found during BO, averaged over 20 trials, with $\pm$1 standard deviation. To ensure fair comparisons despite outlier contamination, we report the uncontaminated mean best value. In the top row (columns 1–4), using EULBO-EI and under asymmetric outliers, RCaGP clearly outperforms all baselines on Lasso DNA 180D, and ranks first or on par with CaGP on Hartmann 6D, Lunar12D, and Rover60D. SVGP and RCSVGP display the same performance. The bottom row (columns 1–4) shows results with Determinantal Point Process (DPP)-based inducing point selection [moss2023inducing], still in the case of asymmetric outliers. RCaGP remains the top performer on DNA and performs similarly to CaGP on other tasks, with RCSVGP slightly ahead on Hartmann. Substituting EI with Thompson Sampling in DPP-BO leads to equal performance across baselines (Figure˜S1).
Next, Figure˜S2 presents results with TurBO, a BO variant tailored to high dimensions [eriksson2019scalable]. RCaGP exhibits strong average performance, and notably, it is the only method that never ranks last in a statistically significant manner. Lastly, as one might not know beforehand whether outliers will affect the trial, column 5 in Figure 4 shows results for EULBO EI on Hartmann and Lunar without outliers. RCaGP maintains a clear advantage on Hartmann, while all baselines perform similarly on Lunar.
Together, these results highlight RCaGP’s versatility and robustness across inducing point allocation strategies (EULBO, DPP), acquisition functions (EI, TS), BO behavior (global BO, trust-region BO), and noise regimes (with or without outliers). Its consistent superiority supports the central claim: combining robustness from RCGP with the uncertainty calibration of CaGP leads to a synergistic improvement, not just an additive one, across a wide range of BO challenges.
@@ 8 Conclusion
Acknowledgments.
This research was supported by the Research Council of Finland (flagship programme: Finnish Center for Artificial Intelligence, FCAI grants 358958, 345604, and 341763), and the UKRI Turing AI World-Leading Researcher Fellowship, EP/W002973/1. We also acknowledge the computational resources provided by the Aalto Science-IT Project.
@@ Acknowledgments.
  • Appendix˜A gives the results of additional experiments mentioned in the main text. These include experiments on UCI regression datasets with uniform and focused outliers (Table S1 and Table S2), BO results for DPP-BO-TS (Figure S1), TuRBO (Figure S2), and dataset-specific results for the ablation study conducted in Section 7.4, where $c$ is varied.
  • Appendix˜B gives the computational complexity of RCaGP.
  • Appendix˜C formally derives the ELBO for RCaGP introduced in Section 3.2.
  • Appendix˜D contains the proof for the robustness property presented in Proposition˜4.1.
  • Appendix˜E proves Proposition˜4.2, showing that RCaGP’s uncertainty estimates capture the worse-case error over all latent functions.
  • Appendix˜F contains additional theoretical results: a proposition showing the convergence in mean of RCaGP in RKHS norm (Section F.2), and a proposition showing the lack of robustness of CaGP (Section F.1).
  • Appendix˜G provides further details regarding the expert-driven prior mean introduced in Section 5 and employed during the ablation study in Section 7.4.
  • Appendix˜H provides additional details about the UCI regression datasets (Section H.1), the test functions employed in BO (Section H.2), and the outlier-contamination protocols studied in this work (Section H.3).
  • Appendix˜I provides tables that report the hyperparmeters used for UCI regression and high-throughput BO experiments (Section I.2) and the description of the computing resources used for our experiments (Section I.1).
@@ Appendix A Additional experiments
The results of the UCI regression experiments in the presence of uniform and focused outliers are presented in Table˜S1 and Table˜S2, respectively. The MAE and NLL metrics indicate that RCaGP consistently outperforms the baseline methods across most datasets. These findings demonstrate the robustness and versatility of RCaGP in handling different types of outliers. We further conducted experiments using the radial basis function (RBF) kernel on both clean and contaminated datasets. While RCaGP achieves competitive MAE and NLL performance compared to RCSVGP on clean data, it substantially outperforms all baselines on contaminated data. However, this performance improvement comes at the cost of increased computational time compared to the baselines.
The results of Bayesian optimization (BO) experiments using the Thompson Sampling (TS) acquisition function (AF) and Trust-Region BO (TuRBO) are presented in Figure˜S1 and Figure˜S2, respectively. In experiments utilizing the TF acquisition function, no model demonstrates a statistically significant performance advantage. Furthermore, the best values identified by these models are generally lower than those obtained with the BO-EI method. We hypothesize that this performance degradation arises from the reliance of the TS approach on posterior sampling, which may reduce the effectiveness of the robustness properties inherent in RCaGP. Subsequent BO experiments using the TuRBO framework show improved performance across all models and tasks—except for the Lunar task. In this case, TuRBO yields lower results for all models. This suggests that the presence of outliers in the Lunar task may impede TuRBO’s ability to effectively explore and identify the global optimum.
The MAE and NLL plots for varying values of $\epsilon$ in the weight function $w$ are shown in Figure˜S3. Across all datasets, RCaGP consistently yields lower MAE and NLL values than RCSVGP, indicating superior predictive performance. Furthermore, both metrics improve when $\epsilon$ is set to 0.1 or 0.15, but begin to deteriorate once $\epsilon$ exceeds 0.2. This trend suggests that aligning $\epsilon$ with the true proportion of outliers in the data enhances model robustness and overall performance.
Table S1: UCI Regression datasets results in the presence of uniform outliers with $\epsilon=0.1$. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits. Bolded results refer to the best baseline. Lower is better.
@@
Table S2: UCI Regression datasets results in the presence of focused outliers with $\epsilon=0.1$. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits. Bolded results refer to the best baseline. Lower is better.
@@
Table S3: UCI Regression datasets results with RBF kernel. Average test set mean absolute error, negative log-likelihood, and clock-time (in seconds), with 1 std, for 20 train-test splits. Bolded results refer to the best baseline. Lower is better.
Asymmetric outliers No outliers
Boston Energy Yacht Parkinsons Boston Energy Yacht Parkinsons
SVGP MAE 1.145 ± 0.052 1.138 ± 0.040 1.235 ± 0.044 1.002 ± 0.100 0.731 ± 0.052 0.908 ± 0.026 0.767 ± 0.063 0.811 ± 0.078
NLL 1.691 ± 0.043 1.736 ± 0.037 1.732 ± 0.041 1.590 ± 0.079 1.433 ± 0.053 1.430 ± 0.023 1.420 ± 0.078 1.448 ± 0.070
clock-time 0.930 ± 0.280 1.170 ± 0.070 0.650 ± 0.060 0.720 ± 0.260 0.940 ± 0.030 1.230 ± 0.150 0.870 ± 0.240 0.670 ± 0.130
CaGP MAE 1.060 ± 0.095 0.971 ± 0.090 1.133 ± 0.092 0.955 ± 0.110 0.488 ± 0.042 0.334 ± 0.024 0.512 ± 0.042 0.674 ± 0.077
NLL 1.611 ± 0.086 1.535 ± 0.071 1.679 ± 0.088 1.556 ± 0.080 1.278 ± 0.040 1.106 ± 0.011 1.263 ± 0.042 1.362 ± 0.067
clock-time 2.470 ± 0.060 3.540 ± 0.080 0.640 ± 0.100 0.870 ± 0.080 2.410 ± 0.060 3.540 ± 0.160 0.770 ± 0.060 0.830 ± 0.100
RCSVGP MAE 0.680 ± 0.122 0.456 ± 0.041 0.733 ± 0.114 0.840 ± 0.187 0.512 ± 0.074 0.558 ± 0.035 0.640 ± 0.074 0.540 ± 0.083
NLL 1.364 ± 0.080 1.200 ± 0.026 1.381 ± 0.084 1.518 ± 0.116 1.302 ± 0.063 1.300 ± 0.034 1.318 ± 0.047 1.417 ± 0.065
clock-time 1.500 ± 0.090 2.560 ± 0.100 0.970 ± 0.090 0.970 ± 0.090 1.530 ± 0.040 2.780 ± 0.120 1.130 ± 0.080 0.720 ± 0.060
RCaGP MAE 0.733 ± 0.156 0.353 ± 0.053 0.651 ± 0.095 0.913 ± 0.198 0.379 ± 0.050 0.326 ± 0.023 0.365 ± 0.035 0.534 ± 0.078
NLL 1.405 ± 0.118 1.173 ± 0.024 1.336 ± 0.061 1.531 ± 0.152 1.241 ± 0.046 1.144 ± 0.007 1.226 ± 0.034 1.307 ± 0.068
clock-time 2.550 ± 0.060 3.590 ± 0.090 0.790 ± 0.060 0.900 ± 0.030 2.470 ± 0.030 3.740 ± 0.170 0.940 ± 0.070 1.070 ± 0.240
Student-t MAE 0.962 ± 0.037 1.015 ± 0.028 0.907 ± 0.061 1.025 ± 0.079 0.896 ± 0.041 0.970 ± 0.025 0.815 ± 0.071 0.965 ± 0.074
NLL 1.619 ± 0.039 1.606 ± 0.024 1.564 ± 0.067 1.658 ± 0.085 1.593 ± 0.063 1.572 ± 0.026 1.559 ± 0.100 1.627 ± 0.092
clock-time 1.030 ± 0.040 1.350 ± 0.070 0.790 ± 0.060 0.730 ± 0.040 1.090 ± 0.030 1.460 ± 0.060 0.920 ± 0.050 0.730 ± 0.010
@@ Appendix C Evidence Lower Bound for RCaGP
$=-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{k}_{i}({\bf{X}},{\bf{X}}){\bf{J}}_{\bf{w}}^{-1/2})-\frac{1}{2}{\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}})^{\top}{\bf{J}}_{\bf{w}}^{-1/2}\hat{\mu}_{i}({\bf{X}}))+\hat{\mu}_{i}({\bf{X}})^{\top}\nu\$
@@
$=\int-\frac{1}{2}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}\,q_{i}({\bf{f}})\,d{\bf{f}}+\int{\bf{f}}^{\top}\nu\,q_{i}({\bf{f}})\,d{\bf{f}}$
$\quad-\int\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}})\,q_{i}({\bf{f}})d{\bf{f}}$ (S8)
$=-\frac{1}{2}\int{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}\,q_{i}({\bf{f}})\,d{\bf{f}}+{\mathbb{E}}_{q_{i}({\bf{f}})}[{\bf{f}}^{\top}\nu]-\frac{1}{2}C({\bf{x}},{\bf{y}},{\sigma_{\mathrm{noise}}^{2}}),$ (S9)
Next, we apply the identity ${\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-2}{\bf{J}}_{\bf{w}}^{-1}{\bf{f}}={\mathrm{tr}}(\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2}{\bf{f}}{\bf{f}}^{\top}\sigma_{\mathrm{noise}}^{-1}{\bf{J}}_{\bf{w}}^{-1/2})$ (see [altamirano2023robust]):
@@ Appendix D Robustness Property
where we define $)+\hat{{\bf{G}}}={\bf{I}}-\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}})(\hat{{\bf{B}}}^{-1/2}({\bf{I}}-\hat{{\bf{S}}}^{+}\hat{{\bf{S}}}$$)^{+}$ and + denotes the Moore-Penrose inverse.
@@
PIF of RCaGP We now prove Proposition˜4.1: Let ${\bf{f}}\sim{\mathcal{GP}}(m,k)$ denote the RCaGP prior, and let $i\in{0,\dots,n}$ represent the number of actions in RCaGP. Assume the observation noise is $\varepsilon\sim{\mathcal{N}}(\mathbf{0},{\sigma_{\mathrm{noise}}^{2}}{\bf{I}})$. Define constants $C_{k}^{\prime}\in;k=1,2,3$, which are independent of $y_{m}^{c}$. For any $i$ and assuming $\sup_{{\bf{x}},y}w({\bf{x}},y)<\infty$, the PIF of RCaGP is given by:
@@
$\leq\left(n\sup_{{\bf{x}},y}\,4\sigma^{-2}_{\mathrm{noise}}\,w^{2}({\bf{x}},y)\,\left(n\sup_{\hat{{\bf{x}}},\hat{y}}\sigma^{4}_{\mathrm{noise}}/2\;w^{-2}(\hat{{\bf{x}}},\hat{y})+\sum_{k=1}^{n}{\bf{K}}_{kk}\right)\right)\bar{C}_{1}\bar{C}_{2}-n=\bar{C}_{3}$ (S61)
Next, we derive the bound for the term (2). Following [altamirano2023robust], we have that
@@
for a constant $q<+\infty$. The second equality comes from the definition of $\tilde{{\bf{v}}}_{i}$, and the fourth row follows the Cauchy-Schwarz inequality. Finally, the last inequality holds since $\tilde{{\bf{K}}}_{i}=(\tilde{{\bf{C}}}_{i}^{-1}-\tilde{{\boldsymbol{\Sigma}}}_{i})$. Applying results from [altamirano2023robust], we obtain
@@
where $\bar{C}_{11}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{7}+\bar{C}_{8})^{2}$ and $\bar{C}_{12}=q\bar{C}_{6}\|{\bf{K}}\|_{F}2(\bar{C}_{9}+\bar{C}_{10})^{2}$. The terms $\bar{C}_{7},\bar{C}_{8},\bar{C}_{9},\bar{C}_{10}$ correspond to $\tilde{C}_{6},\tilde{C}_{8},\tilde{C}_{7},\tilde{C}_{9}$ in [altamirano2023robust].
@@ Appendix E RCaGP Capture the Worst-case Errors
where $\mathcal{D}$} denotes the observation set. Then, applying [kanagawa2018gaussian, Lemma 3.9] provides
@@
since ${\bf{x}}\neq{\bf{x}}_{j}$, it holds that $k^{w}({\bf{x}},{\bf{X}})=k({\bf{x}},{\bf{X}})$. By definition, we have $k^{w}({\bf{X}},{\bf{X}})=\tilde{{\bf{K}}}={\bf{K}}+\frac{{\sigma_{\mathrm{noise}}^{2}}}{2}{\bf{J}}_{\bf{w}}$, and following in [wenger2022posterior, Eq. (S42)], it holds that $\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}=\tilde{{\bf{C}}_{i}}$. Therefore, we obtain
@@
$\qquad\;k^{w}({\bf{x}},{\bf{X}})\tilde{{\bf{C}}}_{i}\tilde{{\bf{K}}}\tilde{{\bf{C}}}_{i}k^{w}({\bf{X}},{\bf{x}})$ (S100)
@@
The proof follows immediately from Proposition˜4.2 and notice that $h/\|h\|_{{\mathcal{H}}_{k^{w}}}$ has unit norm. Corollary˜E.1 implies that the computational uncertainty provides the pointwise bound relative to RCGP’s mathematical posterior, and the combined uncertainty is the pointwise bound relative to the shifted latent function.